OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 3x2 + 5x + 1 = 0

OpenStudy (anonymous):

Delta = b^2 - 4ac x = ( - b + - sqrt Delta ) / 2a a = 3 b = 5 c = 1 D = 25 - 4 * 3 * 1 D = 25 - 12 D = 13 x = (-5 + - 3.6) /6 x = - 5 + 3.6 / 6 x = -1.4/6 x = -14/60 x = - 7/30 x = - 5 - 3.6 / 6 x = -8.6 / 6 x = -86/60 x = -43/30